How is the energy of a photon dependent on the frequency and wavelength? Based on that, how many photons are emitted in a nitrogen gas laser pulse with a wavelength of 337 nm and energy of 3.83 mJ?


1
Expert's answer
2021-01-05T12:03:34-0500

"E = hν = \frac{hc}{λ} \\

λ = 337 \; nm \\

E_{total} = 3.83 \; mJ \\

E = \frac{hc}{λ} \\

= \frac{6.626 \times 10^{-34} \times 3 \times 10^8 }{337 \times 10^{-9}}"

"= 58.9852 \times 10^{-20} \;J" (energy of one photon)

Number of photons:

"N = \frac{E_{total}}{E} \\

= \frac{3.83 \times 10^{-3}}{58.9852 \times 10^{-20}} \\

= 6.4932 \times 10^{15} \;photons"

Answer: "6.4932 \times 10^{15} \;photons"